What is the water solubility of BaCO3 (Ksp = 2.6 x 10-9) in a solution containing 0.60 M CO32-(aq)? Answer: 8.6E-9 Check
Answer #1

[CO32-] = 0.60 M

At equilibrium:

BaCO3 <----> Ba2+ + CO32-

   s 0.6 + s

Ksp = [Ba2+][CO32-]

2.6*10^-9=(s)*(0.6+ s)

Since Ksp is small, s can be ignored as compared to 0.6

Above expression thus becomes:

2.6*10^-9=(s)*(0.6)

2.6*10^-9= (s) * 0.6

s = 4.333*10^-9 M

Answer: 4.3*10^-9 M